OpenStudy (anonymous):

I need to calculate the scientific notation 8.3 x 10e6 kg/2.5 x 10e9 cme3. I can't convert the kg into cm. Anyone know how to do this problem?

OpenStudy (anonymous):

kg doesnt convert to cm, kg is a weight where as cm is a distance does that make sense? is the problem given to you in fraction form? if so you can have your final answer be kg/cm

OpenStudy (anonymous):

Looking at it it will be \[\frac{kg}{cm^3}\]

OpenStudy (anonymous):

The problem can be viewed as: \[\frac{8.3}{2.5}*\frac{10^6}{10^9}*\frac{kg}{cm^3}\]

OpenStudy (anonymous):

cm^3 is a volume so that can be done and \[1cm ^{3} = 1ml\]

OpenStudy (anonymous):

wait a thats a volume not a weight. okay either way there not going to cancel out so you'll multiply the top across and the bottom across and then your lable will just be kg/\[cm ^{3}\]

OpenStudy (anonymous):

That's a density problem (how much a known volume weighs) and kg*cm^-3 is considered a derived unit. \[\left(\begin{matrix}8.3*10^6 \\ 2.5*10^9\end{matrix}\right)kg*cm ^{-3}=3.32*10^{-3} kg*cm ^{-3}\]